# Re-Reading Improves Reasoning in Large Language Models

### v3

@@ 6 Limitations
In this paper, we introduce a simple yet effective prompting method for enhancing reasoning in LLMs and conduct extensive experiments to validate its effectiveness. Despite our best efforts, there may be still some limitations that remain in our study. Our investigation primarily revolves around empirical studies with extensive experiments to validate Re2, similar to most works in prompting research Zheng et al. (2023); Yin et al. (2023); Gao et al. (2023). Future efforts will include more theoretical analyses to provide a solid foundation. Additionally, Re2 marginally increases the input length, potentially leading to a slight reduction in efficiency for longer questions during inference. Furthermore, our paper concentrates exclusively on the impact of Re2 within the reasoning domain. Future work will extend the exploration to additional contexts, such as multi-turn dialogue and multi-modal reasoning.
@@ Appendix B Specific Prompting Methods
Appendix C Attention Analysis
@@
Appendix D Case Study
@@
We also conduct a case study to show the effectiveness of our proposed re-reading prompting over the chain-of-thought. We choose two examples from GSM, and the results generated by ChatGPT are listed in Table  12-13. It is evident that our method can better align the evidence in the question with the corresponding explanation hints. We can observe that CoT+Re2 tends to highlight the important evidences in the question before generating the explanation, for example, “In the morning, she gives 15 cups of feed, and in the afternoon, she gives another 25. So …" in Table 12 and “The bonus is worth half a month’s salary, which is …" in Table 13. This observation is also consistent with the right figure in Figure 4.
Appendix E More Cases
@@
Tables 14-18 provide more examples generated by ChatGPT with CoT and CoT+Re2. We also provide several examples generated by davinci-003 and ChatGPT in the Vanilla prompting (e.g. no instruction) in Tables 19-22. They show that ChatGPT with Vanilla directly generates answer in Coin Filp and Date Understanding dataset (Tables 19-20), but still generates CoT output in some other datasets (Tables 21-22).

### v4

@@ 6 Limitations
In this paper, we introduce a simple yet effective prompting method for enhancing reasoning in LLMs and conduct extensive experiments to validate its effectiveness. Despite our best efforts, there may be still some limitations that remain in our study. Our investigation primarily revolves around empirical studies with extensive experiments to validate Re2, similar to most works in prompting research Zheng et al. (2023); Yin et al. (2023); Gao et al. (2023). Future efforts will include more theoretical analyses to provide a solid foundation. Additionally, Re2 marginally increases the input length, leading to a slight reduction in efficiency for longer questions during inference. Future work will explore more scenarios except reasoning, such as multi-turn dialogue and multi-modal reasoning.
@@ 7 Ethics
Acknowledgements
We would like to thank the anonymous reviewers for their constructive comments and suggestions. This work was supported by the National Natural Science Foundation of China (No.61925203 and No.U22B2021).
@@ Appendix B Specific Prompting Methods
Appendix C GPT-4o-mini Experiments
LLMs are rapidly evolving, with more powerful models emerging frequently. To assess the effectiveness of Re2 on newer, more advanced models, we tested it on GPT-4o-mini, specifically the gpt-4o-mini-2024-07-18 version  [^3] https://platform.openai.com/docs/models/gpt-4o-mini. The results, presented in Figure 8 and Figure 9, demonstrate that Re2 continues to perform effectively on these more advanced LLMs.
Appendix D Attention Analysis
@@
Appendix E Perplexity Analysis
Figure 7: The perplexity of generating the question or the ground-truth answer with increasing reading times.
For the explanation about “overly repeating questions encourages LLMs to repeat the question rather than generate the answer” in Section 3.4, we conducted an experiment. This experiment aims to investigate the likelihood of generating questions versus generating ground-truth responses as reading times of the question increased. We pose two research questions: (1) Does the probability of generating the question as the output increase with more reading times? (2) Does the probability of generating the ground-truth response decrease with more reading times?
Specifically, for each question in the GSM8k dataset, we provide the LLM with the question with varying repetition times as input, and set the LLM’s output as the question itself or its ground-truth response. We then calculate the perplexity of generating both the question and the ground-truth answer. Perplexity serves as an indicator reflecting the likelihood of generating a sequence, with lower perplexity indicating a higher likelihood. These experiments are conducted using the Llama 2. The perplexity results are summarized in Figure 7.
The results reveal two key findings: (1) The perplexity of generating questions decreases with increasing reading times, suggesting that the LLM finds it easier to generate the question. (2) With the exception of when reading times = 2, the perplexity of generating the ground-truth response increases overall. This finding aligns with optimal performance observed when the question is read only twice. Moreover, as reading times increase, the LLM appears to be less inclined to generate the answer.
Appendix F Case Study
@@
We also conduct a case study to show the effectiveness of our proposed re-reading prompting over the chain-of-thought. We choose two examples from GSM, and the results generated by ChatGPT are listed in Table  14-15. It is evident that our method can better align the evidence in the question with the corresponding explanation hints. We can observe that CoT+Re2 tends to highlight the important evidences in the question before generating the explanation, for example, “In the morning, she gives 15 cups of feed, and in the afternoon, she gives another 25. So …" in Table 14 and “The bonus is worth half a month’s salary, which is …" in Table 15. This observation is also consistent with the right figure in Figure 4.
Appendix G More Cases
@@
Tables 16-20 provide more examples generated by ChatGPT with CoT and CoT+Re2. We also provide several examples generated by davinci-003 and ChatGPT in the Vanilla prompting (e.g. no instruction) in Tables 21-24. They show that ChatGPT with Vanilla directly generates answer in Coin Filp and Date Understanding dataset (Tables 21-22), but still generates CoT output in some other datasets (Tables 23-24).
Dataset Type N Answer Example
GSM8K Math 1319 Number Josh decides to try flipping a house. He buys a house for \$80,000 and then puts in \$50,000 in repairs. This increased the value of the house by 150%. How much profit did he make?
SVAMP Math 1000 Number After resting they decided to go for a swim. The depth of the water is 15 times Dean’s height. Dean is 4 feet taller than Ron. If Ron stands at 13 feet. How deep was the water?
ASDIV Math 2096 Number There are 3300 bananas in Janice’s banana collection. Janice also has 5 crayons. If the bananas are organized into 75 groups, how big is each group?
AQUA Math 254 Option The original price of an item is discounted 22%. A customer buys the item at this discounted price using a \$20-off coupon. There is no tax on the item, and this was the only item the customer bought. If the customer paid \$1.90 more than half the original price of the item, what was the original price of the item? Answer Choices: A)\$61, B)\$65, C)\$67.40, D)\$70, E)\$78.20.
MultiArith Math 600 Number For the school bake sale Robin made 42 cupcakes. If she sold 22 of them and then made 39 more, how many cupcakes would she have?
SingleEq Math 508 Number Alyssa spent half of her allowance going to the movies. She washed the family car and earned 8 dollars. What is her weekly allowance if she ended with 12 dollars ?
AddSub Math 395 Number Mike had 34 peaches at his roadside fruit dish . He went to the orchard and picked peaches to stock up . There are now 86 peaches . how many did he pick ?
CSQA CS 1221 Option Where would you find magazines along side many other printed works? Answer Choices: A)doctor, B)bookstore, C)market, D)train station, E)mortuary.
StrategyQA CS 2290 Yes / No Do the anchors on Rede Globo speak Chinese?
ARC-e CS 2376 Option The shape of the moon appears to change regularly during each month. Which of the following best explains why the shape of the moon appears to change? Answer Choices: A)The Earth turns on its axis, B)The Moon turns on its axis, C)The Moon orbits around the Earth, D)Clouds cover the Moon.
ARC-c CS 1172 Option What is a similarity between sound waves and light waves? Answer Choices: A)Both carry energy, B)Both travel in vacuums, C)Both are caused by vibrations, D)Both are traveling at the same speed.
Date Understanding Sym. 369 Date Yesterday was April 30, 2021. What is the date today in MM/DD/YYYY?
Coin Flip Sym. 500 Yes / No A coin is heads up. Breanna flips the coin. Trey does not flip the coin. Omar flips the coin. Patrice does not flip the coin. Is the coin still heads up? Note that "flip" here means "reverse".
Table 10: Details of reasoning benchmarks. Math: arithmetic reasoning. CS: commonsense reasoning. Sym.: symbolic reasoning.